OpenStudy (anonymous):

4a+3b+c=-9 a-3b+2c=27 11a-2b+3c=28

OpenStudy (anonymous):

c=6 b=-5 a=0

OpenStudy (anonymous):

please show me how you got those answers so i can do by myself for next ones

OpenStudy (anonymous):

c=-4a-3b-9 a-3b+2c=27 11a-2b+3c=28 Replace all occurrences of c with the solution solution The value of a variable that makes an equation true. found by solving the last equation equation A mathematical statement that says two expressions have the same value; any number sentence with an = . for c. In this case, the value substituted is -4a-3b-9. c=-4a-3b-9 a-3b+2(-4a-3b-9)=27 11a-2b+3c=28 Replace all occurrences of c with the solution solution The value of a variable that makes an equation true. found by solving the last equation equation A mathematical statement that says two expressions have the same value; any number sentence with an = . for c. In this case, the value substituted is -4a-3b-9. c=-4a-3b-9 a-3b+2(-4a-3b-9)=27 11a-2b+3(-4a-3b-9)=28

OpenStudy (anonymous):

totally confused

OpenStudy (anonymous):

i have had four different people tell me different but yours is the only one that was verified when i calculate the substitutions

OpenStudy (anonymous):

okay, good. So you got it right?

OpenStudy (anonymous):

trying to figure out your equation

OpenStudy (anonymous):

need help with anything else?

OpenStudy (anonymous):

can u plz work that equation step by step so i can understand what it is you did

OpenStudy (anonymous):

its too long to work out. I just did it in my head. Ill write it out tomorrow for you. Im going to go to bed now.

OpenStudy (anonymous):

ok thank you so much for your help

OpenStudy (anonymous):

write above equation in system matrix notation and use cramer rule to solve unknown variables

OpenStudy (anonymous):

Got the steps for you, if you dont understand anything please tell me I kind of rushed it, sry.